OpenStudy (anonymous):

lim as x approaches infinity (1/(k^3 +1))^1/k = 1 but why?

OpenStudy (anonymous):

One method you can do is take the natural log and see if that approaches anything.

OpenStudy (anonymous):

\[ \lim_{k\to \infty}\left(\frac 1{k^3+1}\right)^{1/k} \]

OpenStudy (anonymous):

Considering how \(x = e^{\ln(x)}\) when \(x>0\), we can take the \(\ln\) and see what happens to the expression: \[ \frac 1k \bigg( \ln(1) - \ln(k^3+1)\bigg) = -\frac{\ln(k^3+1)}{k} \]

OpenStudy (anonymous):

Now, typically it is understood that \(x\) dominates \(\ln x\), so this will converge to \(0\). \(e^{\ln x} = e^0 = 1\)

OpenStudy (anonymous):

oh, ok that makes a lot of sense, thanks

OpenStudy (anonymous):

As k tends to infinity direct sustituion of limits will gt u 0^0 which is an undefined form inmath

OpenStudy (anonymous):

@wio Follow this method works fine for some case for more complicated cases there r some theorems u can use to handle these undefined forms